Find all solutions of the equation below. (If there are more answer boxes than answers, enter NONE in the extra boxes.)
x4 - 4096 = 0
x = (smaller real solution)
x = (larger real solution)
x = (smaller imaginary part)
x = (larger imaginary part)
i dont understand! could someone help explain it to me thanks!!!

Found 2 solutions by stanbon, solver91311:
Answer by stanbon(75887)

x^4 - 4096 = 0
x^4 - 2^12 = 0
Factor to get:
(x^2-2^6)(x^2+2^6) = 0
Factor to get:
(x-2^3)(x+2^3)(x+8i)(x-8i) = 0
x = 8 or x = -8 or x = -8i or x = 8i
======================================
Cheers,
Stan H.

Answer by solver91311(24713)





First thing to notice is that you have the difference of two squares:


, so you can factor:



Using the Zero Product Rule:

, or



If , then we can again factor the difference of two squares:

and we have our two real roots, namely 8 and -8.

On the other hand, if , then so which is to say giving us our conjugate pair of complex roots.

In order, your answers should be: